OpenStudy (anonymous):

((4)/(x))-((1)/(x+1))=((??)/(x(x+1)))

OpenStudy (zehanz):

You have to write it as one fraction. The new denominator is the product of x and x+1. To make this happen, multiply both numerator and denominator of the first fraction with x+1. Then multiply num. and den. of the second fraction with x: \[\frac{ 4 }{ x }-\frac{ 1 }{ x+1 }=\frac{ 4(x+1) }{ x(x+1) }-\frac{ x }{ x(x+1) }\]Now write as one fraction and simplify.

OpenStudy (anonymous):

So the answer would be: \[\frac{ 3x+4 }{ x(x+1) }\]?

OpenStudy (zehanz):

Right!